OpenStudy (anonymous):

There are 6 orange, 2 red, 3 white, and 4 green candies in a box. Joe chooses a candy, looks at it,replaces it, and then draws another. What is the probability of him drawing 2 orange candies?

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

6/16 add allup then put the amount of candy there is

OpenStudy (helpsos):

nvm

OpenStudy (openstudier):

Since we are replacing, these are independent events (the probability of the first event doesn't affect the second event), so the formula is: \[P(A and B)=P(A) * P(B)\]

OpenStudy (helpsos):

sorry :(

OpenStudy (anonymous):

I believe it's \(\left(\dfrac{6}{16}\right)^2\)

OpenStudy (anonymous):

Hmmmmm.

OpenStudy (anonymous):

Yup, @OpenStudier is right.

OpenStudy (openstudier):

So what would be the P(A) and P(B)?

pooja195 (pooja195):

6/16 is the answer @OpenStudier yr just explaing it too much

OpenStudy (anonymous):

Oh, ok.

OpenStudy (openstudier):

@pooja195, one does not simply not explain on OpenStudy if @help123please doesn't understand. And also, you are wrong.

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

And yes, I don't understand still.

OpenStudy (openstudier):

Yes. What is the probability of drawing one orange?

OpenStudy (anonymous):

omg. typo

OpenStudy (anonymous):

there are 2 red candies

OpenStudy (anonymous):

sorry bout that :/

OpenStudy (openstudier):

So what is the probability of getting one orange?

OpenStudy (anonymous):

One orange? 1/ 15?

OpenStudy (openstudier):

No. If there is 6 oranges from the total, one orange is 6/15=2/5

OpenStudy (anonymous):

Ohhh. Sorry bout that.

OpenStudy (anonymous):

Okay. But it is asking for 2 oranges.

OpenStudy (anonymous):

2/5*2?

OpenStudy (anonymous):

[which is 8/10?]

OpenStudy (openstudier):

So if we know the probability of getting one orange, what is the proabability of getting two oranges twice in a row (what the problem is asking)?

OpenStudy (anonymous):

Is that right? 8/10?

OpenStudy (openstudier):

No, look back up at my earlier explanation.

OpenStudy (anonymous):

Oh. Okay...Lets see.

OpenStudy (anonymous):

What is A and B?

OpenStudy (anonymous):

What do they represent?

OpenStudy (anonymous):

@Kitt020912

OpenStudy (openstudier):

A and B is what I refered to the probability of events A and B.

OpenStudy (anonymous):

Okay.

OpenStudy (openstudier):

So the answer would be: (2/5) x (2/5)... Understand?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Que pensez-vous, Kitt?

OpenStudy (anonymous):

so 4/25?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks.:)